OpenStudy (anonymous):

For the functions f(x)=x^5 and g(x)=x+4, find the following: f(g(1)): Did I do this one right? f(1+4)=f(5)=5^5=3,125 g(f(1)): What about this one? g(1^5)=g(1)=1+4=5 f(g(x)): f(x+4)=(x+4)^5 did I start that right? :/ and if so, is that as simplified as it gets? g(f(x)): g(x^5)=(x^5) + 4 is that right? and is that as simplified as it can get? f(t)g(t): this one i'm a bit confused on :/ would it just be (x^5)(x+4)=x^6+4x^5 ? Thank you!!! :)

OpenStudy (anonymous):

g(1)=1+4=5 f(5)=5^5=3125

OpenStudy (anonymous):

f(t)g(t)=(t^5)(t+4)=t^6+4t^5 not x

OpenStudy (agent0smith):

(x^5) + 4 you can drop the brackets, that's it (x+4)^5 doesn't get any more simplified unless you like doing really boring unnecessary work.

OpenStudy (anonymous):

f(1)=1^5=1 g(1)=1+4=5 g(x)=x+4 f(x+4)=(x+4)^5 = x^5+20 x^4+160 x^3+640 x^2+1280 x+1024

OpenStudy (agent0smith):

^yeah there is your unnecessary work. Don't do that.

OpenStudy (anonymous):

oh right haha forgot... wait so would these be correct then? f(g(1))=3,125 g(f(1))=5 f(g(x))=(x+4)^5 g(f(x))= x^5+4 f(t)g(t)=t^6+4t^5 ?? are those the final correct answers then? :)

OpenStudy (anonymous):

haha okay @agent0smith :) i won't :)

OpenStudy (anonymous):

thanks!!! :)